OpenStudy (anonymous):

HOW DO I FIGURE THIS OUT \[6\sqrt{2} - 3\sqrt{2}\]

OpenStudy (kinggeorge):

\[6\sqrt{2} - 3\sqrt2 = \sqrt{2} \cdot (6-3) =3\sqrt2\]

OpenStudy (anonymous):

how do i do these? \[\sqrt{90} , \sqrt{125}, 4\sqrt{20}\]

OpenStudy (kinggeorge):

What do you want to do with them? Just simplify them as much as possible?

OpenStudy (anonymous):

no have an anser like \[9\sqrt{10} or 10\sqrt{9} examples\]

OpenStudy (anonymous):

This one i dont get either \[5\sqrt{10} * 3\sqrt{?}\]

OpenStudy (anonymous):

i ment to put * \[3\sqrt{8}\]

OpenStudy (kinggeorge):

Let's start with the first one.\[\sqrt{90} = \sqrt{9} \cdot \sqrt{10} = 3\sqrt{10}\]similarly, \[\sqrt{125} = \sqrt{25} \cdot \sqrt{5} = 5\sqrt5\]and\[4\sqrt{20} = 4\sqrt{4} \cdot \sqrt5 = 4 \cdot 2 \cdot \sqrt5 = 8\sqrt5\]

OpenStudy (anonymous):

\[7\sqrt{72}\]

OpenStudy (kinggeorge):

As for that last problem, \[5\sqrt{10} \cdot 3\sqrt8 = 15 \cdot \sqrt{10} \cdot \sqrt8 = 15 \cdot \sqrt{80}=15 \cdot \sqrt{16} \cdot \sqrt5 = 60\sqrt5\]Is it clear what you have to do for these problems now?

OpenStudy (anonymous):

i dont get the one i posted or the 4/20 one.

OpenStudy (anonymous):

would this one \[\sqrt{8} * \sqrt{3}\] = \[\sqrt{24}\]?

OpenStudy (anonymous):

would this one \[\sqrt{8} * \sqrt{3}\] = \[\sqrt{24}\]?

OpenStudy (kinggeorge):

That would be correct. Good job! As for the other problems, give me a few more seconds to finish typing up the response...

OpenStudy (anonymous):

KK :)

OpenStudy (kinggeorge):

Well, both the 7/72 and 4/20 are basically the same problem with different numbers. You have an expression with some number outside the radical (4 and 7), and a number inside (20 and 72). At the beginning, you can just ignore the number outside. Then simplify the radical as much as possible. In this case, \[\sqrt{72} = \sqrt{36} \cdot \sqrt2 = 6\sqrt2\]and \[\sqrt{20} = \sqrt4 \cdot \sqrt5 =2\sqrt5\]Then, to finish it all off, you multiply by the number you originally had in front. So,\[7\sqrt{72}=7\cdot(6\sqrt2) = 42\sqrt2\]\[4\sqrt{20}=4\cdot (2\sqrt5)=8\sqrt5\]And those would be the final answers. Also, with the square root of 24, it would be better to then simplify like so\[\sqrt{24} = \sqrt4 \cdot \sqrt6 = 2\sqrt6\]

OpenStudy (anonymous):

i see... ok. Thanks so much!

OpenStudy (kinggeorge):

Could you now simplify \[4\sqrt{45}\]just for a little bit more practice?